two points of charges +2 columb and +6 columb repel each other with afor of 12 neton of charge columb is given to each of these charges will be the force noe
 

Grade:12

1 Answers

ASHISH KUMAR
2 years ago
given :-  q1 = 2  ,  q2 = 6  & F = 12 N
 
F = \frac{K q1 q2 }{r^{^{2}}}
12 = k (2)(6)/r^2
12 = k 12/r^2
k/r^2 = 1
now when -4 Coulomb charge given to each charge then new charges will be
 
q1' = 2-4 = -2       &      q2 ' = 6-4 = 2
 
New force = Fnew = k q1' q2' / r^2= 1 * (-2) *(2) = -4 N 
or  Fnew = 4N ( attractive)